Aerotech International Inc. of Montreal has supported a research and development (R&D) department that for many years has been the sole contributor to the company’s new products. The R&D activity is an overhead cost centre that provides services only to in-house manufacturing departments (four different product lines), all of which produce aerospace-related products.

The department has never sold its services outside, but because of its long history of success, larger manufacturers of aerospace products have approached Aerotech to hire its R&D department for special projects. Because the costs of operating the R&D department have been spiralling uncontrollably, Aerotech’s management is considering taking on these outside contracts to absorb the increasing costs. However, managers don’t have any cost basis for charging R&D services to outsiders, and they need to gain control of their R&D costs. Management decides to implement an activity-based costing system in order to determine the charges for both outsiders and the in-house users of the department’s services.

R&D activities fall into four pools with the following annual costs:
Market analysis $1,015,000 Product development $3,459,960
Product design 2,260,800 Prototype testing 1,340,160

Analysis determines that the appropriate cost drivers and their usage for the four activities are as follows:
Activities Cost Drivers Total Estimated Drivers
Market analysis Hours of analysis 14,500 hours
Product design Number of designs 2,400 designs
Product development Number of products 90 products
Prototype testing Number of tests 480 tests

Collapse question part

(a)

Calculate the activity-based overhead rate for each activity cost pool.
Market analysis $

per hour of analysis
Product design $

per design
Product development $

per product
Prototype testing $

per test
Answer #1
Activities Activity cost Cost Drivers Total Estimated Drivers Overhead rate
Market analysis $1,015,000 Hours of analysis 14,500 hours 1,015,000/14,500 = $70 per hour of analysis
Product design 2,260,800 Number of designs 2,400 designs 2,260,800/2,400 = $942 per design
Product development $3,459,960 Number of products 90 products 3,459,960/90 = $38,444 per product
Prototype testing 1,340,160 Number of tests 480 tests 1,340,160/480 = $2,792 per test

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.